BexleyBrooks:

Need some more help on Math. 1. 5+4(n+9)=-3 what is N? 2. 5(k-2)-8k=-34 what is K? 3. 2=11+3(m+3) what is M? 4. -2(p-5)+7p=-5 what is P? 5. 4a-2a(a+9)=6 what is A? 6. 7-4(d-3)=23 what is D? 7. 8x-11(x-2)=-8 what is X?

HotPockets14:

give me 5 minutes

Jayisbetter:

Let's solve each equation: 5 + 4 ( � + 9 ) = − 3 5+4(n+9)=−3 Distribute the 4: 5 + 4 � + 36 = − 3 5+4n+36=−3 Combine like terms: 4 � + 41 = − 3 4n+41=−3 Subtract 41 from both sides: 4 � = − 44 4n=−44 Divide by 4: � = − 11 n=−11 5 ( � − 2 ) − 8 � = − 34 5(k−2)−8k=−34 Distribute the 5: 5 � − 10 − 8 � = − 34 5k−10−8k=−34 Combine like terms: − 3 � − 10 = − 34 −3k−10=−34 Add 10 to both sides: − 3 � = − 24 −3k=−24 Divide by -3: � = 8 k=8 2 = 11 + 3 ( � + 3 ) 2=11+3(m+3) Distribute the 3: 2 = 11 + 3 � + 9 2=11+3m+9 Combine like terms: 2 = 20 + 3 � 2=20+3m Subtract 20 from both sides: 3 � = − 18 3m=−18 Divide by 3: � = − 6 m=−6 − 2 ( � − 5 ) + 7 � = − 5 −2(p−5)+7p=−5 Distribute the -2: − 2 � + 10 + 7 � = − 5 −2p+10+7p=−5 Combine like terms: 5 � + 10 = − 5 5p+10=−5 Subtract 10 from both sides: 5 � = − 15 5p=−15 Divide by 5: � = − 3 p=−3 4 � − 2 � ( � + 9 ) = 6 4a−2a(a+9)=6 Distribute the -2a: 4 � − 2 � 2 − 18 � = 6 4a−2a 2 −18a=6 Combine like terms: − 2 � 2 − 14 � = 6 −2a 2 −14a=6 Subtract 6 from both sides: − 2 � 2 − 14 � − 6 = 0 −2a 2 −14a−6=0 This is a quadratic equation. You can solve for � a using the quadratic formula. 7 − 4 ( � − 3 ) = 23 7−4(d−3)=23 Distribute the -4: 7 − 4 � + 12 = 23 7−4d+12=23 Combine like terms: − 4 � + 19 = 23 −4d+19=23 Subtract 19 from both sides: − 4 � = 4 −4d=4 Divide by -4: � = − 1 d=−1 8 � − 11 ( � − 2 ) = − 8 8x−11(x−2)=−8 Distribute the -11: 8 � − 11 � + 22 = − 8 8x−11x+22=−8 Combine like terms: − 3 � + 22 = − 8 −3x+22=−8 Subtract 22 from both sides: − 3 � = − 30 −3x=−30 Divide by -3: � = 10 x=10

luhivqqcherry:

..

luhivqqcherry:

she looked that up

Jayisbetter:

@luhivqqcherry wrote:
she looked that up
Im not a girl

xXChudLord420Xx:

@jayisbetter why did you format it like that

Jayisbetter:

@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom

luhivqqcherry:

@jayisbetter wrote:
@luhivqqcherry wrote:
she looked that up
Im not a girl
ok

xXChudLord420Xx:

@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
damn

HotPockets14:

1). N = -11 2). k= 8 3). m = -6 4). p = 5/9 5). A = 7+37/2 6). D = -1 7). x = 10

luhivqqcherry:

@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .

xXChudLord420Xx:

@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
shut up nerd

BexleyBrooks:

@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
no fighting on my post! I will report you

luhivqqcherry:

@xxchudlord420xx wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
shut up nerd
imagine

Jayisbetter:

@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
yeah be quiet

luhivqqcherry:

@jayisbetter wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
yeah be quiet
no

xXChudLord420Xx:

@luhivqqcherry wrote:
@jayisbetter wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
yeah be quiet
no
the first stage of greif is denial

luhivqqcherry:

@xxchudlord420xx wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
yeah be quiet
no
the first stage of greif is denial
ok

BexleyBrooks:

@xxchudlord420xx wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@luhivqqcherry wrote:
@jayisbetter wrote:
@xxchudlord420xx wrote:
@jayisbetter why did you format it like that
because your mom
cringe .
yeah be quiet
no
the first stage of greif is denial
both of you shut up! I will not tolerate arguing

Illuminous:

Copy and pasted..

Illuminous:

You can tell, sorry x.x

BexleyBrooks:

@illuminous wrote:
You can tell, sorry x.x
what is copy and pasted?